OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9

OpenStudy (immanuelv):

you could do subtraction

OpenStudy (anonymous):

lel r u srs m9? x - 2y - 14 = 0 x - 14 = 2y (x - 14) / 2 = y x/2 - 7 = y

OpenStudy (anonymous):

x + 3y = 9 x + 3y - 9 = 0 x - 9 = -3y (x - 9)/-3 = y x/-3 + 3 = y

OpenStudy (anonymous):

(1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

those are the choics

OpenStudy (anonymous):

(x-2y) - (x+3y) = 14-9 -5y = 5

OpenStudy (anonymous):

The first one is (12, -1)

OpenStudy (anonymous):

what do i do with -5y=5

OpenStudy (anonymous):

???

OpenStudy (anonymous):

wot m8

OpenStudy (anonymous):

what will be the value of y if -5y =5

OpenStudy (anonymous):

y = -1

OpenStudy (anonymous):

im really lost

OpenStudy (anonymous):

to simplify the equation, work pemdas backwards. Just skip to division for -5y = 5. Now divide both sides by -5. y = -1

OpenStudy (anonymous):

is it -1,-12

OpenStudy (anonymous):

\[(x-2y) - (x+3y) = 14-9\] \[-5y = 5\] y = -1 x - 2y = 14 x - 2(-1) = 14 x + 2 = 14 so x will be?

OpenStudy (anonymous):

The first problem you asked about is (12, -1)

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

Actually, they're both (12, -1)

OpenStudy (anonymous):

so the answer is (12,-1)

OpenStudy (anonymous):

Indeed

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

anytime